OpenStudy (anonymous):

How would I determine whether the following function is odd, even, or neither without sketching the graph? f(x)=3x^4-2x^2

OpenStudy (anonymous):

3x^4=f(x)+2x^2 f(x)=x^2(3x^2-2)